OpenStudy (anonymous):

Solve the following system of equations. x + 4y + z = –10 3x – 3y + 6z = –21 x + 2y + 2z = –10

OpenStudy (anonymous):

x + 4y + z = –10 . . . (eq 1) 3x – 3y + 6z = –21 . .(eq 2) x + 2y + 2z = –10 . . (eq 3) divide left and right of the second (eq 2) by 3 3x - 3y + 6z = -21 x - y +2z = -7 subtract this equation from the third x + 2y + 2z = -10 x - y + 2z = -7 ---------------- - 3y = -3 y = -1 enter this value in (eq 1) and (eq 3) x + 4(-1) + z = –10 x + z = -6 . . . . . (eq 4) x + 2(-1) + 2z = –10 x + 2z = –8 . . . . (eq 5) subtract (eq 4) from (eq 5) x + 2z = –8 x + z = -6 ---------- - z = -2 enter this value in (eq 4) x -2 = -6 x = -4 Answer: x = -4 y = -1 z = -2